OpenStudy (anonymous):

Find the linear approximation of the data that passes through the points (5, 3) and (20, 6) y = 15x + 9 y = 15x + 3 y = 1/5x - 1 y = 1/5x + 2

OpenStudy (hybrik):

To find the slope use \[(y2-y1/x2-x1)\]

OpenStudy (hybrik):

(6-3/20-5) = (3/15) = (1/5) is the slope

OpenStudy (hybrik):

so A and B is eliminated

OpenStudy (anonymous):

i think i had this question be4

OpenStudy (anonymous):

i think c

OpenStudy (hybrik):

Then plug in points

OpenStudy (hybrik):

3=1/5(5)+b, 3=1+b, b= 2

OpenStudy (hybrik):

So the correct answer is D not C